OpenStudy (anonymous):

You roll two dice. Let event A be "The first die shows a 2, a 3, or a 5" and event B be "The second die shows an odd number." What is P(A|B)? A. 1/3 B. 1/2 C. 5/6 D. 1/6

OpenStudy (turingtest):

\[P(A|B)={P(A\cap B)\over P(B)}\]what is the \(P(A\cap B)\), which is the probability that both events occur? what is \(P(B)\), the probability that the second die shows an odd number?

OpenStudy (anonymous):

I have no idea

OpenStudy (turingtest):

take into consideration whether these are independent events as well, that will dramatically simplify your calculation

OpenStudy (turingtest):

does rolling one die have an effect on the other?

OpenStudy (anonymous):

No

OpenStudy (turingtest):

so then, by the definition of independence, \(P(A\cap B)=P(A)P(B)\)so then what does the expression for \(P(A|B)\) become?

OpenStudy (anonymous):

total no. of exhaustive events=6*6=36 favourable events are {21,23,25,31,33,35,51,53,55} no. of favourable events=9 P=? i don't know where i am wrong.